OpenStudy (anonymous):

simplify 5^4√7-8^3√7-2^4√7-^3√7

OpenStudy (agreene):

\[5^4\sqrt{7-8^3}\sqrt{7-2^4}\sqrt{7-1^3}\sqrt 7\] Is that the question? it's hard to interpret what you mean without parenthetical/bracketed statements.

OpenStudy (anonymous):

no. its exactly how it looks .

OpenStudy (agreene):

\[5^4\sqrt 7 -8^3\sqrt 7 -2^4\sqrt7 -1^3\sqrt7=96\sqrt7\]

OpenStudy (anonymous):

thats what i got but these are the answers: 3^4√7-9^3√7 -6√14 -6√6 3^4√14-9^3√14

OpenStudy (agreene):

O_o none of those are even close... they are all negative. and are around ~1,000.001 ish in decimal...

OpenStudy (anonymous):

i donnnno

OpenStudy (agreene):

me either :X

OpenStudy (anonymous):

well the last part is -^3√7 not -1^3√7 ? if that makes a diff

OpenStudy (anonymous):

The 1st answer is OK. Need to rite expression differently:\[5\sqrt[4]{7}-8\sqrt[3]{7}-2\sqrt[4]{7}-\sqrt[3]{7}\]\[\sqrt[4]{7}(5-2)-\sqrt[3]{7}(8+1)\]\[\sqrt[4]{7}*3-\sqrt[3]{7}*9\]\[3\sqrt[4]{7}-9\sqrt[3]{7}\]